# Lamb-shift-induced switching of energy transfer in open quantum batteries

### 21 Mar 2026 v1

@@ I INTRODUCTION
Our results reveal a clear manifestation of mode-selective energy transfer governed by the Lamb-shift-modified resonance condition. We demonstrate that by tuning the driving frequency to the renormalized eigenfrequencies $\lambda_{\pm}$, one can selectively optimize energy distribution between the charger and the battery-a phenomenon we call the ”switching effect.” Furthermore, we propose a frequency correction strategy based on the Lamb-shifted eigenvalues to restore ergotropy that would otherwise degrade due to environmental interference. These findings offer actionable insights for designing controllable quantum energy devices in interference-prone environments.
The remainder of this paper is organized as follows: In Sec.I, we describe the model setup and derive the master equation with Lamb-shift corrections. In Sec.II, we present the numerical results and the physical analysis of the switchable charging dynamics. Finally, the experimental feasibility and conclusions are summarized in Sec. III and Sec. IV.
@@ II Model and Theoretical Framework
where $\gamma_{a}$ is the decay rate and $\mathcal{D}_{\hat{c}}[\hat{\rho}]=\hat{c}\hat{\rho}\hat{c}^{\dagger}-\frac{1}{2}\{\hat{c}^{\dagger}\hat{c},\hat{\rho}\}$ denotes the standard dissipator. The mean thermal excitation number is $N=[\exp(\omega/k_{B}T)-1]^{-1}$. Crucially, we incorporate the Lamb shift $\Delta_{L}$, which originates from the principal value contribution of the system-bath interaction under the Born-Markov and secular approximations.. Although often neglected in the weak-coupling limit, $\Delta_{L}$ represents a non-negligible frequency renormalization in realistic open systems, ensuring a physically comprehensive description of the energy-level structure [11].
@@
where $\mathcal{D}[\hat{O}]\langle\hat{A}\rangle=\langle\hat{O}^{\dagger}\hat{A}\hat{O}-\frac{1}{2}\{\hat{O}^{\dagger}\hat{O},\hat{A}\}\rangle$. In the zero-temperature limit ($N=0$), the dynamics of the annihilation operators reduce to a set of coupled first-order differential equations [15]:

### 02 Apr 2026 v2

@@ I INTRODUCTION
Our results reveal a clear manifestation of mode-selective energy transfer governed by the Lamb-shift-modified resonance condition. We demonstrate that by tuning the driving frequency to the renormalized eigenfrequencies $\lambda_{\pm}$, one can selectively optimize energy distribution between the charger and the battery-a phenomenon we call the “switching effect.” Furthermore, we propose a frequency correction strategy based on the Lamb-shifted eigenvalues to restore ergotropy that would otherwise degrade due to environmental interference. These findings offer actionable insights for designing controllable quantum energy devices in interference-prone environments.
The remainder of this paper is organized as follows: In Sec.II, we describe the model setup and derive the master equation with Lamb-shift corrections. In Sec.III, we present the numerical results and the physical analysis of the switchable charging dynamics. Finally, the experimental feasibility and conclusions are summarized in Sec. IV and Sec. V.
@@ II Model and Theoretical Framework
where $\gamma_{a}$ is the decay rate and $\mathcal{D}_{\hat{c}}[\hat{\rho}]=\hat{c}\hat{\rho}\hat{c}^{\dagger}-\frac{1}{2}\{\hat{c}^{\dagger}\hat{c},\hat{\rho}\}$ denotes the standard dissipator. The mean thermal excitation number is $N(t)=[\exp(\omega/k_{B}T)-1]^{-1}$. Crucially, we incorporate the Lamb shift $\Delta_{L}$, which originates from the principal value contribution of the system-bath interaction under the Born-Markov and secular approximations.. Although often neglected in the weak-coupling limit, $\Delta_{L}$ represents a non-negligible frequency renormalization in realistic open systems, ensuring a physically comprehensive description of the energy-level structure [11].
@@
where $\mathcal{D}[\hat{O}]\langle\hat{A}\rangle=\langle\hat{O}^{\dagger}\hat{A}\hat{O}-\frac{1}{2}\{\hat{O}^{\dagger}\hat{O},\hat{A}\}\rangle$. In the zero-temperature limit ($N(t)=0$), the dynamics of the annihilation operators reduce to a set of coupled first-order differential equations [15]: